OpenStudy (anonymous):

Need geometry help. The equation of a circle is (x + 6)2 + (y + 2)2 = 16. The point (-6, 2) is on the circle. What is the equation of the line that is tangent to the circle at (-6, 2)?

OpenStudy (anonymous):

I suppose you mean \((x+ 6)^2 + (y+2)^2 = 16\), correct?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Wouldn't you have to take the derivative to find the slope, then plug the points (-6,2) and your slope into the form y= mx+b to find the y intercept? Then rewrite the equation for the tangent line?

OpenStudy (anonymous):

I hope you know where you're heading Jhannybean because this is not single variable analysis anymore.

OpenStudy (jhannybean):

Oh hahaha i was just guessing!! nevermind then :P

OpenStudy (anonymous):

Line perpendicular to the line connecting the point and the center that also goes through the given point?

OpenStudy (jhannybean):

Yesss

OpenStudy (anonymous):

The slope is zero at the point so it will just be the line y = 2. But this is only via visual inspection actually proving this is more difficult.

OpenStudy (anonymous):

ok cool thanks for the help

OpenStudy (jhannybean):

 omg this is so not to scale :|

OpenStudy (jhannybean):

maybe that's how it's drawn? :\ not sure...

OpenStudy (jhannybean):

Probably not, haha, i tried!!

OpenStudy (anonymous):

thanks for trying

OpenStudy (anonymous):

Visit the wolfram link for the graph he was trying to draw.

OpenStudy (jhannybean):

Ohhhhh I see it now!!!

OpenStudy (anonymous):

thanks again for helping

OpenStudy (anonymous):

I'm not sure how you can prove it formally without multivariable calculus.

OpenStudy (jhannybean):

Haha.i shall be learning that next semester! :P

OpenStudy (anonymous):

Obviously, the picture is THE thing!

OpenStudy (anonymous):

How to prove it with analysis? Notice that the circle is a manifold without boundary. We can define the chart \(\alpha(\theta) = (4\cos(\theta) - 6, 4\sin(\theta) - 2)\) \). Then just consider the tangent space of the chart at the angle corresponding to that point which is going to be \(\pi/2\)

OpenStudy (anonymous):

Oh my!

OpenStudy (jhannybean):

@Alchemista are you done with all math?past multivariable calc? just wondering.

OpenStudy (anonymous):

One is never done with all math. INFINITY DON'T YOU KNOW?

OpenStudy (jhannybean):

lol